Express the equation y=2x+3 in the standard form and find two solutions. Is (2,3) its solution?
Answer by srinivas.g(540)

standard form of an equation of line is ax+by =c
given equation is y =2x+3
add -2x on both sides
y -2x = 2x+3 -2x
y-2x= 3
-2x+y =3
standard form : -2x+y =3
given point (x,y)= (2,3)
-2(2)+3= -4+3
=-1
but -2x+y=3
hence the point (2,3) is not the solution